# Benchmarking Vision Language Models for Cultural Understanding

### v1

@@ 4 Dataset Analysis
Figure 3: Comparative analysis of data by country. The figure presents three aspects: (A) unique counts of images, questions, and answers, (B) average lengths of questions and answers, and (C) average number of answers and inter-annotator agreement scores across countries, showcasing variations and trends in CulturalVQA.
@@
According to the pie chart in Fig. 4, food-related questions are most prevalent, accounting for 31.6% of the dataset, followed closely by traditions and rituals, which represent 28.6% and 22.6% respectively. Thus, roughly 50% of the questions in our dataset probe for cultural understanding of the intangible aspects of culture (rituals and traditions)! The word clouds generated from the collected answers in Fig. 4 reveal diverse expressions of rituals and traditions represented by terms like hamam (Turkey) and meskel (Ethiopia). Further, the food category includes diverse items such as feijoada (Brazil), fufu (Nigeria), and vada (India) indicating a geo-diverse culinary scope. While the clothing category is the least prevalent in the dataset, it shows the highest variety in terms of collected answers. The drink category is notably one of the smallest, both in terms of the size and number of unique answers.
@@ 5 Benchmarking VLMs on CulturalVQA
Since we sourced questions from different annotator groups across countries, it is imperative to ask if the disparity in model performance across countries is due to the disparity in inherent difficulty levels of questions across countries. To investigate this, we analyze the correlation between the model performance and the average question length (see Fig. 3 (B) for average question length across countries). We use average question length as a proxy for question difficulty because we believe shorter questions probe more direct knowledge, whereas longer questions require more nuanced, detailed cultural knowledge. We find the correlation between question length and model performance to be approx. zero (ranging between -0.3 and 0.3) for most models, with the exception of GPT-4, which exhibits a moderate negative Pearson correlation of -0.51. This indicates that shorter questions are easier to answer than longer ones. However, the variation in average question length (ranging from 8 to 12 words) is too small to make a conclusive claim.
Another factor potentially affecting the disparity in model performance is the variable number of human answers per question across countries (see Fig. 3 (C)). These human answers are used as the reference answers in the LAVE metric making it more rigid for countries with fewer references and vice-versa. To investigate this, we compute the correlation between model performance and the average number of answers per question. We find a very low Pearson correlation varying in either direction depending on the model (from -0.3 to 0.2), indicating that the disparity in the number of human answers does not meaningfully affect the disparity in model performance across countries.
@@ 9 Acknowledgements
We would like to extend our gratitude to David Ifeoluwa Adelani for connecting us with Masakhane, Fırat Öncel for assisting with annotators in Turkey, Saba Ahamadi for securing annotators from Iran, and Qian Yang for sourcing annotators from China. We also appreciate the valuable feedback provided by Ibrahim Alabdulmohsin on the early draft. The technical support from the Mila IDT team in managing the computational infrastructure is greatly appreciated. Additionally, Aishwarya Agrawal received support from the Canada CIFAR AI Chair award throughout this project. Karolina Stańczak was supported by the Mila P2v5 grant and the Mila-Samsung grant. This project was generously funded by a research grant from Google.

### v2

@@ 4 Dataset Analysis
Figure 3: Comparative analysis of data by country. The figure presents three aspects: (A) unique counts of images, questions, and answers, (B) average lengths of questions and answers, and (C) average number of answers per question and inter-annotator agreement scores across countries, showcasing variations and trends in CulturalVQA.
@@
According to the pie chart in Fig. 4, food-related questions are most prevalent, accounting for 37.3% of the dataset, followed closely by traditions and rituals, which represent 26.1% and 18% respectively. Thus, roughly 44% of the questions in our dataset probe for cultural understanding of the intangible aspects of culture (rituals and traditions)! The word clouds generated from the collected answers in Fig. 4 reveal diverse expressions of rituals and traditions represented by terms like hamam (Turkey) and meskel (Ethiopia). Further, the food category includes diverse items such as feijoada (Brazil), fufu (Nigeria), and vada (India) indicating a geo-diverse culinary scope. While the clothing category is the least prevalent in the dataset, it shows the highest variety in terms of collected answers. The drink category is notably one of the smallest, both in terms of the size and number of unique answers.
@@ 5 Benchmarking VLMs on CulturalVQA
We also calculate the Pearson correlation between human and model performance across countries. For open-source models, we observe a relatively low correlation, ranging from 0.1 to 0.4. Interestingly, for closed-source models like Gemini and GPT-4, we find a stronger correlation of 0.69 and 0.75, respectively. This suggests that the factors affecting human performance similarly influence the performance of these closed-source models. However, from Fig. 8, when comparing human and model performance using the same metric, we find that closed-source models still lag behind humans for all countries indicating that while these models follow human performance trends, there is still a marked gap in their cultural understanding compared to humans. This gap is even more pronounced for open-source models, which show an even larger discrepancy in model performance compared to human performance across all countries.
@@
Figure 11: Scatter plot of GPT-4V performance versus average question length across different countries
@@
Since we sourced questions from different annotator groups across countries, it is imperative to ask if the disparity in model performance across countries is due to the disparity in inherent difficulty levels of questions across countries. To investigate this, we analyze the correlation between the model performance and the average question length (see Fig. 3 (B) for average question length across countries). We use average question length as a proxy for question difficulty because we believe shorter questions probe more direct knowledge, whereas longer questions require more nuanced, detailed cultural knowledge. As the variance in absolute question lengths is small, looking at ranks is more meaningful, which is why we use Spearman for this analysis. We find a weak Spearman rank correlation between question length and model performance ranging between -0.3 and 0.3 for most models, with the exception of GPT-4 and Gemini, which exhibit a moderate negative correlation of -0.52 on average. Further, Fig. 11 presents a scatter plot of GPT-4 performance vs. average question lengths across countries. We can see that for most countries except Brazil, Canada, and the USA, the variance in question lengths is small. This suggests that question length is not a significant factor behind the disparity in model performance across countries.
Another factor potentially affecting the disparity in model performance is the variable number of human answers per question across countries (see Fig. 3 (C)). These human answers are used as the reference answers in the LAVE metric making it more rigid for countries with fewer references and vice-versa. To investigate this, we compute the correlation between model performance and the average number of answers per question across countries. We again use Spearman correlation since the variance in absolute values of average numbers of answers per question across countries is small. We find a very low correlation ranging between -0.3 and 0 across models, indicating that the disparity in the number of human answers does not meaningfully affect the disparity in model performance across countries.
@@ 9 Acknowledgements
We would like to extend our gratitude to David Ifeoluwa Adelani for connecting us with Masakhane, Fırat Öncel for assisting with annotators in Turkey, Saba Ahamadi for securing annotators from Iran, and Qian Yang for sourcing annotators from China. We also appreciate the valuable feedback provided by Ibrahim Alabdulmohsin on the early draft. The technical support from the Mila IDT team in managing the computational infrastructure is greatly appreciated. We would also like to thank Chris Emezue for his assistance in helping with payments for the annotators. Additionally, Aishwarya Agrawal received support from the Canada CIFAR AI Chair award throughout this project. Karolina Stańczak was supported by the Mila P2v5 grant and the Mila-Samsung grant. This project was generously funded by a research grant from Google.